# Enhancing Table Recognition with Vision LLMs: A Benchmark and Neighbor-Guided Toolchain Reasoner

### v2

@@ 3.4.1. Experimental Setup
Table Border Quality. Table borders indicate structural information and are critical for parsing tasks. We evaluate the impact of border visibility and completeness on table recognition performance by considering scenarios including unclear table borders and missing borders. Additionally, we explore the effect of border changes in the table by thickened table borders.
@@
Geometric Deformation. Geometric deformation caused by viewing angles or operations can disrupt the geometric consistency of tables, which is common in visual table parsing, especially when tables are scanned or photographed. We evaluated the robustness of VLLMs against geometric deformations by testing them under tilt conditions of 20° and excessive tilt conditions of 40°.
3.4.2. Discussion and Analysis
@@
Figure 4 shows the detailed experimental results. We observe that blurring and overexposure can lead to loss of text information in table images and inaccurate fine-grained text recognition, thus negatively impacting the performance of table content recognition. In contrast, underexposure has little effect on table recognition performance. Moreover, VLLM demonstrates limited efficacy when processing skewed tables, resulting in a substantial decrease in accuracy under such conditions.
@@
Although table borders are often considered essential for conveying structured information, modifications such as fading or removing these borders have minimal impact on the performance of VLLMs in TR tasks. This result suggests that VLLMs do not heavily rely on borders in TR. VLLMs only pay slight attention to the structural information the borders provide when thickening. Structural information provided by the borders.
@@
Figure 4. Evaluation results of bottleneck scenarios: abbreviations UE (Underexposure), OE (Overexposure), MB (Missing Borders), UB (Unclear Borders), TB (Thickened Borders).
@@
Figure 5. Illustration of a pipeline for table image preprocessing leveraging a toolkit of lightweight vision models.
4. Neighbor-Guided Toolchain Reasoner
@@
Through our in-depth analysis of VLLMs’ performance on benchmark tasks and bottleneck analysis, we have identified that the quality of the input image emerges as a critical factor that constrains the performance of these VLLMs. Consequently, improving input image quality is essential for enhancing VLLMs’ capability to recognize and interpret structured image data more effectively. To address this bottleneck, we propose the NGTR framework.
4.1. Model Overview
@@
NGTR enhances input image quality by applying various tool combinations tailored to scenarios such as low resolution, overexposure, and noise interference, among others. To select a suitable combination of tools, we employ a similarity-based neighbor retrieval module to obtain samples that resemble each input instance. These samples guide the generation of multiple tool invocation plans. Subsequently, the tool invocation experience learning module executes each toolchain plan and generates the corresponding structured data based on the neighbor samples. It evaluates the effectiveness of different plans to select a relatively better plan. In the final processing stage, we utilize a reflection-driven tool selection module. This module integrates iterative tool invocation and dynamic feedback to refine the processing flow. The optimized image resulting from this process is subsequently input into the VLLMs, which utilize their powerful reasoning capabilities to generate the structured data representation of the table.
Table 3. Specific descriptions of built tools in the toolkit.
@@
Binarization The binarization tool converts images to black and white, highlighting key features for easier extraction.
@@
4.2. Toolkit Preparation
4.3. Similarity-based Neighbor Retrieval
Neighbor retrieval methods enable VLLMs to retrieve similar neighbor samples, providing richer contextual information. In the NGTR framework, we hypothesize that images with similar features exhibit similar results after being processed by the same image preprocessing toolchain. Consequently, the processing results of neighbor samples could guide the selection of a potentially optimal toolchain for test samples. We first retrieve images that are similar to the target task images from a sample set (a subset of the training data containing 100 samples). Subsequently, we employ a prompting template to leverage the VLLM’s planning capabilities to generate tool invocation plans. The retrieval process can be formally described as follows:
(1) $\text{Retrieval}(I^{\text{test}},\mathcal{D}^{\prime},f)=\arg\max\left[f\left(I^{\text{test}},I^{i}\right)\right]_{i=1}^{|\mathcal{D}^{\prime}|},$
where $I^{\text{test}}$ represents an image from the test set, $\mathcal{D}^{\prime}$ denotes a subset of the training dataset, and $f$ is the similarity measurement function. In this paper, we combine the ORB (Oriented FAST and Rotated BRIEF) algorithm with the Hamming distance as $f$ to measure the similarity between images. Specifically, the feature vectors are first extracted using the ORB algorithm for the task image and a candidate example image. Subsequently, the Hamming distance between the feature vectors of the two images is calculated as a measure of similarity. Then, we guide the VLLMs to generate multiple tool invocation plans for the image. The generation process can be formally expressed as follows:
(2) $\text{VLLM}(\mathcal{T},\mathcal{N}(I^{\text{test}}))\rightarrow\{p_{1},p_{2},\ldots,p_{n}\},$
4.4. Tool Invocation Experience Learning
@@
(3) $I_{p_{i}}=f_{p_{i}}(I),\quad i\in\{1,2,\ldots,n\},\vspace{-0.04in}$
where $f_{p_{i}}$ denotes the image preprocessing process corresponding to the tool invocation plan $p_{i}$. Next, we use a simple prompt template to instruct VLLMs to generate an markup sequence.
(4) $\hat{H}_{P_{i}}=\text{VLLM}(I_{P_{i}},P),$
where $P$ represents the prompt to instruct VLLMs generation.
Subsequently, we evaluate the prediction results based on the example labels. The evaluation process employs the tree edit distance-based similarity (TEDS) metrics to quantify the accuracy of the VLLMs output. The specific details of TEDS can be found in the appendix E. By following this process, we calculate a quantitative score for each toolchain, enabling selection of a suitable plan.
4.5. Reflection-driven Tool Selection
Let $I^{(t-1)}$ denote the image before the $t$-th operation and $I^{(t)}$ denote the image after the $t$-th operation. The VLLMs computes a decision value $\gamma^{(t)}$ to determine whether to accept the operation:
(5) $\gamma^{(t)}=\text{reflect}(I^{(t-1)},I^{(t)}),$
This step-by-step module enhances the interaction between the VLLMs and the target image, ensuring the accuracy of the final task outcome. More importantly, introducing this module enables downstream researchers and developers to flexibly customize and expand the toolkit without worrying about the impact of poorly performing expanded tools on the final results, thereby significantly improving the versatility and transferability of our framework. In the last step, we use a simple prompt template to instruct VLLM to generate a markup sequence and obtain the result of table recognition.
5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
In this study, we utilize three widely-used table recognition datasets: SciTSR (chi2019complicated, ), PubTabNet (zhong2020image, ), and WTW (long2021parsing, ), each offering unique characteristics and challenges, as shown in Table 4. SciTSR is a dataset comprising tables extracted from the scientific literature, and the image quality in this dataset is relatively high. In contrast, the image resolution of PubTabNet is 72 pixels per inch, and its overall image quality is relatively low. WTW contains images collected from the wild, introducing a variety of extreme cases, such as tilt, blur, and table curvature, which significantly increase the task’s difficulty. These datasets encompass diverse table types as well as various unique visual challenges, providing a robust foundation for benchmarking in this study.
Table 4. Characteristics of selected table recognition datasets.
Dataset Tables Source
Markup Sequence
Annotations
Textual
Annotations
Year
SciTSR 15K ArXiv $\times$ $\checkmark$ 2019
PubTabNet 568K PubMed $\checkmark$ $\checkmark$ 2020
WTW 14.6K Wild Scenarios $\times$ $\times$ 2021
Table 5. Performance comparison of methods on the SciTSR, PubTabNet, and WTW datasets. ”-” indicates the method’s lack of results (specific reasons are provided in the implementation details). Best scores in the lightweight OCR model category are highlighted in blue, while best scores in the prompt tuning in VLLMs category are highlighted in green.
@@ 5.1.1. Datasets
5.1.2. Baselines
We select six Vision Large Language Models (VLLMs), including Phi, Llama, GPT-mini, Qwen, GPT, and Gemini, as baseline models for comparison. These models’ selection and implementation details are described in appendix D. Additionally, we select three representative deep learning-based methods as baselines for comparison: EDD (zhong2020image, ) based on sequence modeling, LGPMA (qiao2021lgpma, ) based on cell bounding box detection, and LORE (xing2023lore, ) based on cell point center detection.
5.1.3. Evaluation Metrics
As for the evaluation metrics of TR, we use a similarity metric based on Tree-Edit Distance (TEDS) (zhong2020image, ) and the structure-aware TEDS-Struct metric. The specific details of TEDS can be found in the appendix E. We employ two evaluation metrics for the hierarchical tasks described in Section 3.2.1: accuracy (ACC) and micro-averaged F1 score (F1-score). Specifically, ACC is used to evaluate cell-level tasks (excluding merged cell detection) and table-level tasks; the F1-score is utilized for row- and column-level tasks and merged cell detection.
5.1.4. Implementation Details
For the PubTabNet dataset, since the test set labels are not public, we choose to use the validation set for evaluation. We randomly select 1,500 images from the validation set of PubTabNet as research samples. The experimental results show that the EDD and LGPMA baseline model’s performance on this sample is consistent with its initial reported results, which shows that our sampling is well-representative. For the SciTSR and WTW datasets, we use their complete test sets for evaluation. Since the WTW dataset does not provide content information for table recognition, we do not report its TEDS scores.
For LORE, since it is mainly aimed at table structure recognition but not table content recognition, we only report its performance scores for table structure recognition. As for EDD, since its model training requires a large amount of end-to-end annotated data, and SciTSR and WTW lack corresponding labeled data, its performance on these datasets has not been evaluated. For LGPMA, since the model relies on table content for training and the WTW dataset does not provide table content labels, we did not evaluate its performance on the WTW dataset.
5.2. Main Results Analysis
@@
5.2.1. Performance Analysis of NGTR
@@
As shown in Table 5, the experimental results compare our NGTR framework with baseline methods. We use GPT and Gemini, which performs well in benchmark tests, as the framework backbone. The main results show that our framework achieves significant performance improvements on the PubTabNet dataset, mainly attributed to our framework’s enhanced VLLMs robustness when dealing with low-quality inputs. On the SciTSR dataset, our framework also outperforms all VLLMs baselines, further verifying our framework’s effectiveness.
@@
In the WTW dataset, our method and all VLLM-based baselines demonstrate suboptimal performance compared to traditional lightweight baseline methods. However, despite these challenges, our approach significantly enhances the effectiveness of VLLMs, suggesting that our framework imparts a degree of robustness to VLLMs in complex environments. Nevertheless, it highlights persistent challenges that remain difficult to address in VLLMs. Section 5.2.4 presents a more detailed analysis of VLLMs performance on the WTW dataset.
5.2.2. Open-source and Proprietary VLLMs
@@
The advanced open-source VLLMs demonstrate capabilities comparable to proprietary VLLMs in this task. As a representative of open-source VLLMs, Llama exhibits outstanding performance, particularly in the table structure evaluation task (TEDS-Struct) on the PubTabNet dataset. Llama achieves a relatively better result, surpassing GPT by 2.73 points and Gemini by 2.36 points. These results confirm the exceptional performance of Llama but also underscore the potential of open-source VLLMs for further research.
5.2.3. Model Scale
We observe that lightweight VLLMs face significant performance limitations in specific tasks. For instance, Phi, a lightweight VLLM with a parameter size of 4.2B, demonstrates consistently low performance across all evaluations. While another lightweight VLLM, GPT-mini, achieves commendable results on the simpler SciTSR dataset, its performance declined significantly on the more challenging PubTabNet dataset, dropping 33% and 21% in TEDS and TEDS-Struct metrics, respectively.
We attribute this phenomenon to two key factors. First, the TR task requires VLLMs to possess fine-grained detection capabilities and the ability to accurately recognize and extract structured information—a difficult challenge. Due to their relatively small parameter sizes and limited computational power, lightweight VLLMs struggle to achieve the required recognition accuracy and input robustness. Second, the TR task demands the generation of accurate markup sequence representations of tables, imposing stringent requirements on the VLLMs’ generation capabilities. We observe that lightweight VLLMs often fail to produce parsable markup sequences after multiple attempts, mainly when tasked with complex scenarios such as large, intricate tables or ambiguous image inputs. This limitation significantly reduces their final performance scores, a phenomenon not observed in larger-scale VLLMs.
5.2.4. WTW Dataset: A Challenge for VLLMs
@@
Table 5 presents the experimental results on the WTW dataset. These results indicate that prompt tuning based VLLMs methods still exhibit significant gaps compared to traditional lightweight OCR methods, highlighting the challenges VLLMs face when processing datasets with wild scenarios. An in-depth analysis of the experimental findings reveals that the WTW dataset primarily emphasizes table structure information. Performance declines notably when processing tables with numerous empty cells, unevenly distributed text, skewed, rotated, or densely packed text. Further analysis of the outputs suggests that VLLMs tend to ignore cells lacking semantic content. While this behavior helps avoid processing irrelevant data, it also limits their ability to effectively capture the structural information of tables with many blank cells.
In conjunction with the analysis of table borders in Section 3.4, we speculate that VLLMs primarily rely on cell content and its logical position in the image when completing the table recognition task, with relatively limited use of table border information. This dependence restricts the performance of VLLMs in recognizing complex table structures to some extent.
@@
Table 6. Ablation study results on key components of the proposed framework. ”EXP” denotes the tool invocation experience learning module, while ”REF” represents the reflection-driven tool selection module.
@@
5.3. Ablation Study w.r.t Key Components
The results of our ablation experiments on the NGTR framework are shown in Table 6. We randomly select 1,000 images from the SciTSR dataset for ablation studies. We ablate two key components: tool invocation experience learning and reflection-driven tool selection module. Note that we do not ablate the toolkit, as it is the foundational module of our framework and cannot be removed. The results highlight the critical contributions of these components to the overall performance of the NGTR framework.
@@
Effectiveness of Tool Invocation Experience Learning. To evaluate the effectiveness of VLLMs in optimizing toolchains using neighbor samples, we performed an ablation study by removing the tool invocation experience learning module. In this experiment, we only prompted VLLMs to generate a tool invocation plan and applied it directly to the test samples. Without the tool invocation experience learning module, the experimental results showed that while the VLLMs was able to generate a valid tool invocation plan, the lack of effective validation of the generated plan led to a significant performance decline. This further validates the critical role of the tool invocation experience learning module in improving the NGTR framework performance.
@@
Effectiveness of Reflection-driven Tool Selection. To evaluate the effectiveness of introducing the reflection-driven tool selection module for stepwise backtracking validation in the final execution stage, we ablate this module. We apply all the tools in the toolchain to the task images in a single pass, using them as input for the generation stage. Experimental results show that, without the reflection module for stepwise backtracking validation, VLLMs cannot effectively supervise the processing procedure, which may lead to the incorporation of unsuitable tools for the current sample, thereby affecting performance. This performance drop is mainly reflected in the reduction of the TEDS indicator, which may be due to the loss of text details caused by the incorrect noise reduction and binarization application.
Figure 6. Impact of tool count (left) and multi-path count (right) on performance.
5.4. Tool Invocation Analysis
To further investigate the NGTR framework’s preferences in tool invocation, we calculated the tool usage rates for each test dataset (i.e., the proportion of samples that invoke a particular tool among all samples where tools are invoked). The results are presented in Table 7. As shown, there are notable differences in the frequency of tool usage rates across different datasets.

### v3

@@ 3.4.1 Experimental Setup
Table Border Quality. Table borders indicate structural information in table elements. We evaluate the impact of border visibility and completeness on table recognition performance by considering scenarios including unclear table borders and missing borders. Additionally, we explore the effect of border changes in the table by thickened table borders.
@@
Geometric Deformation. Geometric deformation caused by viewing angles or operations can disrupt the geometric consistency of tables. We evaluated the robustness of VLLMs against geometric deformations by testing them under tilt conditions of 20° and excessive tilt conditions of 40°.
3.4.2 Discussion and Analysis
@@
Figure 4 shows the detailed experimental results. We observe that blurring and overexposure can lead to loss of text information in table images and inaccurate fine-grained text recognition, thus negatively impacting the performance of table content recognition. Moreover, VLLM demonstrates limited efficacy when processing skewed tables, resulting in a substantial decrease in accuracy under such conditions.
@@
Although table borders are often considered essential for conveying structured information, fading or removing these borders has minimal impact on the performance of VLLMs. This result suggests that VLLMs do not heavily rely on borders. VLLMs only pay slight attention to the structural information the borders provide when thickening.
@@
Figure 4: Evaluation results of bottleneck scenarios: abbreviations UE (Underexposure), OE (Overexposure), MB (Missing Borders), UB (Unclear Borders), TB (Thickened Borders).
@@
Figure 5: Illustration of a pipeline for table image preprocessing leveraging a toolkit of lightweight vision models.
4 Neighbor-Guided Toolchain Reasoner
@@
Through our in-depth analysis of VLLMs’ performance on the benchmark, we have identified that improving input image quality is essential for enhancing VLLMs’ capability to recognize and interpret structured image data more effectively. To address this, we propose the NGTR framework.
4.1 Framework Overview
@@
NGTR enhances input image quality by applying various tool combinations tailored to low resolution, overexposure, and noise interference. As shown in Figure 5, we design a similarity-based neighbor retrieval module to select a suitable combination of tools. Subsequently, the tool invocation experience learning module executes each plan and generates the corresponding structured data to evaluate the effectiveness of different plans. Finally, we propose a reflection-driven tool selection module to integrate iterative tool invocation and dynamic feedback to refine the processing flow. The optimized image is then input into the VLLMs, which utilize their powerful reasoning capabilities to generate structured data.
4.2 Toolkit Preparation
4.3 Similarity-based Neighbor Retrieval
Neighbor retrieval methods enable VLLMs to retrieve similar neighbor samples, providing richer contextual information. In the NGTR framework, we hypothesize that images with similar features exhibit similar results after being processed by the same image preprocessing toolchain. Consequently, the processing results of neighbor samples could guide the selection of a potentially optimal toolchain for test samples. We first retrieve images that are similar to the target task images from a sample set. Subsequently, we employ a prompting template to leverage the VLLM’s planning capabilities to generate tool invocation plans. The retrieval process can be formally described as follows:
$\text{Retrieval}(I^{\text{test}},\mathcal{D}^{\prime},f)=\arg\max\left[f\left(I^{\text{test}},I^{i}\right)\right]_{i=1}^{|\mathcal{D}^{\prime}|},$ (1)
where $I^{\text{test}}$ represents an image from the test set, $\mathcal{D}^{\prime}$ denotes a subset of the training dataset, and $f$ is the similarity measurement function. In this paper, we combine the ORB (Oriented FAST and Rotated BRIEF) algorithm with the Hamming distance as $f$ to measure the similarity between images. Then, we guide the VLLMs to generate multiple tool invocation plans for the image. The generation process can be formally expressed as follows:
$\text{VLLM}(\mathcal{T},\mathcal{N}(I^{\text{test}}))\rightarrow\{p_{1},p_{2},\ldots,p_{n}\},$ (2)
Table 3: Specific descriptions of built tools in the toolkit.
@@ 4.3 Similarity-based Neighbor Retrieval
Binarization This tool converts images to black and white, highlighting key features for easier extraction.
@@
4.4 Tool Invocation Experience Learning
@@
$I_{p_{i}}=f_{p_{i}}(I),\quad i\in\{1,2,\ldots,n\},$ (3)
where $f_{p_{i}}$ denotes the image preprocessing tools. Next, we employ a prompt template to guide the VLLMs in generating a markup sequence. Subsequently, we evaluate the prediction results based on the example labels. The evaluation process employs the tree edit distance-based similarity (TEDS) metrics to quantify the accuracy of the VLLMs output. By following this process, we calculate a quantitative score for each toolchain, enabling the selection of a suitable plan.
Table 4: Performance comparison of methods on the SciTSR, PubTabNet, and WTW datasets. ”-” indicates the method’s lack of results (specific reasons are provided in the implementation details). Best scores in the lightweight OCR model category are highlighted in blue, while best scores in the prompt tuning in VLLMs category are highlighted in green.
@@
4.5 Reflection-driven Tool Selection
Let $I^{(t-1)}$ denote the image before the $t$-th operation and $I^{(t)}$ denote the image after the $t$-th operation. The VLLMs computes $\gamma^{(t)}$ to determine whether to accept the operation:
$\gamma^{(t)}=\text{reflect}(I^{(t-1)},I^{(t)}),$ (4)
This step-by-step module enhances the interaction between the VLLMs and the target image, ensuring the accuracy of the final task outcome. More importantly, introducing this module enables downstream researchers and developers to flexibly customize and expand the toolkit without worrying about the impact of poorly performing expanded tools on the final results, thereby significantly improving the versatility and transferability of our framework. In the last step, we design a simple prompt template to instruct VLLM to generate a markup sequence and obtain the result of table recognition.
5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
In this study, we utilize three widely-used table recognition datasets: SciTSR Chi et al. (2019), PubTabNet Zhong et al. (2020), and WTW Long et al. (2021), each offering unique characteristics and challenges. SciTSR is a dataset comprising tables extracted from the scientific literature, and the image quality in this dataset is relatively high. In contrast, the image resolution of PubTabNet is 72 pixels per inch, and its overall image quality is relatively low. WTW contains images collected from the wild, introducing a variety of extreme cases, such as tilt, blur, and table curvature. These datasets encompass diverse table types and various unique visual challenges, providing a robust foundation for benchmarking.
5.1.2 Baselines
We select six Vision Large Language Models (VLLMs), including Phi [^2] https://azure.microsoft.com/en-us/products/phi/, Llama [^3] https://www.llama.com/, GPT-mini, Qwen [^4] https://qwenlm.github.io/blog/qwen-vl/, GPT [^5] https://openai.com/index/hello-gpt-4o/, and Gemini [^6] https://deepmind.google/technologies/gemini/pro/, as baseline models for comparison. Additionally, we select three representative deep learning-based methods as baselines for comparison: EDD Zhong et al. (2020) based on sequence modeling, LGPMA Qiao et al. (2021) based on cell bounding box detection, and LORE Xing et al. (2023) based on cell point center detection.
5.1.3 Evaluation Metrics
As for the evaluation metrics of TR, we use a similarity metric based on Tree-Edit Distance (TEDS) Zhong et al. (2020) and the TEDS-Struct metric. We employ two evaluation metrics for the hierarchical tasks described in Section 3.2.1: accuracy (ACC) and micro-averaged F1 score (F1-score). Specifically, ACC is used to evaluate cell-level tasks (excluding merged cell detection) and table-level tasks; the F1-score is utilized for row- and column-level tasks and merged cell detection.
5.1.4 Implementation Details
For the PubTabNet dataset, We randomly select 1,500 images from the validation set. For the SciTSR and WTW datasets, we use their complete test sets for evaluation. Since the WTW dataset does not provide content information for table recognition, we do not report its TEDS scores.
For LORE, since it is mainly aimed at table structure recognition but not table content recognition, we only report its performance scores for table structure recognition. As for EDD, since its model training requires a large amount of end-to-end annotated data, and SciTSR and WTW lack corresponding labeled data, its performance on these datasets has not been evaluated. LGPMA depends on table content for training, but since WTW lacks content labels, its performance on this dataset was not assessed.
5.2 Main Results Analysis
@@
5.2.1 Performance Analysis of NGTR
@@
As shown in Table 4, the experimental results compare our NGTR framework with baseline methods. The main results show that our framework achieves significant performance improvements on the PubTabNet dataset, mainly attributed to our framework’s enhanced VLLMs robustness when dealing with low-quality inputs. On the SciTSR dataset, our framework also outperforms all VLLMs baselines, further verifying our framework’s effectiveness.
@@
In the WTW dataset, our method and all VLLM-based baselines demonstrate suboptimal performance compared to traditional lightweight baseline methods. However, our approach significantly enhances the effectiveness of these models, suggesting that our framework imparts a degree of robustness in complex environments. Despite this improvement, it also highlights persistent challenges that remain difficult to address. Section 5.2.3 presents a more detailed analysis of performance on the WTW dataset.
5.2.2 Open-source and Closed-source VLLMs
@@
The advanced open-source VLLMs demonstrate capabilities comparable to closed-source VLLMs in this task. As a representative of open-source models, Llama exhibits outstanding performance, particularly in the table structure recognition task (TEDS-Struct) on the PubTabNet dataset. Llama achieves a relatively better result, surpassing GPT by 2.73 points and Gemini by 2.36 points. These results confirm the potential of open-source VLLMs for further research.
5.2.3 WTW Dataset: A Challenge for VLLMs
@@
Table 4 presents the experimental results on the WTW dataset. These results indicate that prompt-tuning based VLLMs methods still exhibit significant gaps compared to traditional lightweight OCR methods, highlighting the challenges VLLMs face when processing datasets with wild scenarios. Further analysis of the outputs suggests that performance declines notably when VLLMs process tables with numerous empty cells, unevenly distributed text, and skewed, rotated, or densely packed text. Further analysis of the outputs suggests that VLLMs tend to ignore cells lacking semantic content. While this behavior helps avoid processing irrelevant data, it also limits their ability to effectively capture the structural information of tables with many blank cells.
@@
Table 5: Ablation study results on key components of the framework. ”EXP” denotes the tool invocation experience learning module, ”REF” represents the reflection-driven tool selection module.
@@
5.3 Ablation Study w.r.t Key Components
@@
Effectiveness of Tool Invocation Experience Learning. We performed an ablation study by removing the tool invocation experience learning module. In this experiment, we only led the VLLMs to generate a tool invocation plan and applied it directly to the test samples. As shown in Table 5, while the VLLMs could generate a valid tool invocation plan, the lack of effective validation of the generated plan led to a significant performance decline. This further validates the critical role of the tool invocation experience learning module in improving the NGTR framework performance.
@@
Effectiveness of Reflection-driven Tool Selection. We performed an ablation study by removing the reflection-driven tool selection module. We apply all the tools to the task images in a single pass. The results are presented in Table 5, without the reflection module for stepwise backtracking validation, VLLMs cannot effectively supervise the processing procedure, which may lead to the incorporation of unsuitable tools for the current sample, thereby affecting performance.
5.4 Tool Invocation Analysis